OpenStudy (anonymous):

Simplify y 2 + 11y - 6y + y 2. y4+ 5 y4+ 5y 2y2+ 5 2y2+ 5y

OpenStudy (anonymous):

add the like terms first

OpenStudy (uri):

y4+11y-6y y4+5y

jhonyy9 (jhonyy9):

@uri are you sure ? y^2 +y^2 = y^4 this is right sure ?

jhonyy9 (jhonyy9):

yes but above you have wrote it different why ?

OpenStudy (uri):

Isn'y y^2+y^2=y^4